### v1

@@ Appendix A.1 Proofs of the Main Results
$\lim_{T\rightarrow\infty}\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p-\varepsilon}^{u}\right)=\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}\tilde{M}_{i}\leq\overline{cv}_{1-p-\varepsilon}^{u}\right)\geq\mathrm{P}\left(\max_{i\leq d}\tilde{M}_{i}\leq\overline{cv}_{1-p-\varepsilon}^{u}\right),$ (A.15)
where the inequality holds because $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i}\leq\max_{i\leq d}\tilde{M}_{i}$. This, combined with (A.14) and the definition of $\overline{cv}_{1-p-\varepsilon}^{u}$ shows that
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\right)\geq 1-p-\varepsilon-o(1).$ (A.16)
Collecting the results from (A.12) and (A.16), we obtain:
$\mathrm{P}\left(\mathrm{MSS}=\overline{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\bigcap\min_{i\in\mathrm{MSS}}M_{i,T}>\overline{cv}_{1-p,T}^{b}\right)$ (A.17)
$\geq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\right)+\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>\overline{cv}_{1-p,T}^{b}\right)-1$
$\geq$ $1-p-\varepsilon-o(1).$
Since $\varepsilon$ is arbitrary, the claim of the lemma follows from (A.17). $Q.E.D.$
Proof of Theorem  2. Since $cv_{1-p,T}^{b}\leq\overline{cv}_{1-p,T}^{b}$ by construction, from Lemma A4(i) and the similar arguments for showing (A.11), we obtain
$\min_{i\in\mathrm{MSS}}M_{i,T}-cv_{1-p,T}^{b}$$\geq$ $(C^{-1}-o_{p}(1))T^{1/2}\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)^{1/2}-O_{p}(1)$
$\geq$ $(C^{-1}-o_{p}(1))\gamma_{T}^{-1/2},$
where the last inequality is by (2.19). Hence
$\mathrm{P}\left(\mathrm{MSS}\subset\widehat{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)$ (A.18)
$\geq$ $\mathrm{P}\left((C^{-1}-o_{p}(1))\gamma_{T}^{-1/2}>0\right)\geq 1-o(1),$
which shows the first claim of the lemma. [^12] This also shows that the finite sample adjustment in Step 7 of the algorithm in Subsection 2.2 is asymptotically negligible.
To show the second claim, we first note that by Lemma A6,
$cv_{1-p,T}^{b}=cv_{1-p,T}^{b}(\overline{\mathrm{MSS}}_{\gamma_{T}})=cv_{1-p,T}^{b}(\mathrm{MSS})$ (A.19)
wpa1. Therefore, by Lemma A5,
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)$$\leq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}(\mathrm{MSS})\right)+o(1)$ (A.20)
$\leq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p+\delta_{T}}^{u}(\mathrm{MSS})\right)+o(1),$
where $\delta_{T}=o(1)$ is a positive sequence. From (A.20), it follows that for any $\varepsilon\in(0,p)$, we have
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p+\varepsilon}^{u}(\mathrm{MSS})\right)+o(1),$
for sufficiently large $T$, which together with Lemma A3 shows that
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq 1-p+\varepsilon+o(1).$ (A.21)
Therefore,
$\mathrm{P}\left(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\bigcap\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)$ (A.22)
$\leq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq 1-p+\varepsilon+o(1).$
To derive the lower bound for $\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p})$, we apply (A.19) to obtain
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\geq\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}(\mathrm{MSS})\right)-o(1).$ (A.23)
@@
This, along with Lemma A5, shows that for any $\varepsilon^{\prime}\in(0,1-p)$
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)$$\geq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p-\delta_{T}}^{u}(\mathrm{MSS})\right)-o(1)$ (A.24)
$\geq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p-\varepsilon^{\prime}}^{u}(\mathrm{MSS})\right)-o(1),$
for sufficiently large $T$. By Lemma A3, we have
$\lim_{T\rightarrow\infty}\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p-\varepsilon^{\prime}}^{u}(\mathrm{MSS})\right)=\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}\tilde{M}_{i}\leq cv_{1-p-\varepsilon^{\prime}}^{u}(\mathrm{MSS})\right).$ (A.25)
From (A.24), (A.25) and the definition of $cv_{1-p-\varepsilon^{\prime}}^{u}(\mathrm{MSS})$, it follows that
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\geq 1-p-\varepsilon^{\prime}-o(1),$
which together with (A.18) shows that
@@
$\mathrm{P}\left(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\bigcap\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)$ (A.26)
$\geq$ $\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq cv_{1-p,T}^{b}\right)+\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)-1$
$\geq$ $1-p-\varepsilon^{\prime}-o(1).$
@@
Since $\varepsilon$ and $\varepsilon^{\prime}$ are arbitrary, the second claim of the lemma follows from (A.22) and (A.26). $Q.E.D.$
Appendix A.2 Auxillary Lemmas
@@
This section presents some auxiliary results that are utilized in proving Lemma 1, Theorems 1 and 2 in the previous section. For any $i=1,\ldots,d$, let $\mu_{i}\equiv$ $\mathrm{E}[R_{i}]$, $\mu_{-i}\equiv$ $\mathrm{E}[R_{-i}]$, $\sigma_{i}^{2}\equiv\mathrm{Var}(R_{i})$ and $\Sigma_{-i}\equiv\mathrm{Var}(R_{-i})$. Define $\tilde{\sigma}_{i}^{2}\equiv\sigma_{i}^{2}-\Gamma_{i,-i}\Sigma_{-i}^{-1}\Gamma_{i,-i}^{\top}$, where $\Gamma_{i,-i}\equiv\mathrm{Cov}(R_{i},R_{-i})$.
@@
where $\tilde{M}_{i}\equiv\left\|(\ell_{d,i}^{\top}\otimes\mathrm{diag}(s_{i,1}^{-1},s_{i,2}^{-1}))\Omega_{d}^{1/2}\mathcal{N}_{d}\right\|_{\infty}$ and $\Omega_{d}=\lim_{T\rightarrow\infty}\Omega_{d,T}$.
For the $t$-th observation $R_{t}^{b}$ in the bootstrap sample $\{R_{t}^{b}\}_{t\leq T_{B}}$, let $R_{-i,t}^{b}$ denotes its subvector excluding $R_{i,t}^{b}$. Define $\tilde{R}_{-i,t}^{b}\equiv(1,(R_{-i,t}^{b})^{\top})^{\top}$ and $\varepsilon_{i,t}^{b}\equiv R_{i,t}^{b}-\theta_{i}^{\top}\tilde{R}_{-i,t}^{b}$. For any proper subset $\mathcal{S}$of $\mathcal{M}$, including the empty set, let $cv_{1-p}^{u}(\mathcal{S})$ represent the ($1-p$)th quantile of $\max_{i\notin\mathcal{S}}\tilde{M}_{i}$, and $cv_{1-p,T}^{b}(\mathcal{S})$ represent the ($1-p$)th quantile of $\max_{i\notin\mathcal{S}}M_{i,T}^{b}$. Moreover, we define:
@@
Let $\overline{cv}_{1-p,T}^{b}\equiv cv_{1-p,T}^{b}(\varnothing)$. Under Assumptions 1 and 2, we have (i) $\overline{cv}_{1-p,T}^{b}=O_{p}(1)$ for any $p\in(0,1)$; and (ii) $\overline{cv}_{1-\gamma_{T},T}^{b}=O_{p}(\gamma_{T}^{-1/2})$.
@@
Under Assumptions 1 and 2, we have $\mathrm{MSS}=\overline{\mathrm{MSS}}_{\gamma_{T}}$ wpa1.
Appendix A.3 Additional Simulation Results
Figure 4: Empirical Distribution of the Intersection Size Between the Estimated and True MSS
Note: This figure illustrates the empirical distribution of the size of the intersection between the estimated and true MSS for four different settings. The upper panel corresponds to the settings of $(K,N,T)=(1,7,120)$ and $(K,N,T)=(1,7,240)$, respectively, while the lower panel corresponds to the settings of $(K,N,T)=(7,1,120)$ and $(K,N,T)=(7,1,240)$. The significance level is set to 0.05.

### v3

@@ Appendix A.1 Proofs of the Main Results
$\lim_{T\rightarrow\infty}\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p+\varepsilon}^{u}(\mathrm{MSS}^{c})\right)=\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}\tilde{M}_{i}\leq cv_{1-p+\varepsilon}^{u}(\mathrm{MSS}^{c})\right).$ (A.17)
@@
This, along with (A.14), (A.16) and the definition of $cv_{1-p-\varepsilon}^{u}(\mathrm{MSS}^{c})$ shows that
$\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq 1-p+\varepsilon+o(1),$ (A.18)
from which, we derive an upper bound for the probability $\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}$:
@@
$\underset{T\rightarrow\infty}{\lim\sup}\mathrm{P}\left(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}\right)$ $=$ $\underset{T\rightarrow\infty}{\lim\sup}\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p,T}^{b}\bigcap\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)$ (A.19)
$\leq$ $\underset{T\rightarrow\infty}{\lim\sup}\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq 1-p+\varepsilon.$
On the other hand, we can obtain a lower bound for the probability $\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}$ as follows:
$\mathrm{P}\left(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}\right)$ $=$ $\mathrm{P}\left(\mathrm{MSS}\subseteq\widehat{\mathrm{MSS}}_{p}\bigcap\widehat{\mathrm{MSS}}_{p}\subseteq\mathrm{MSS}\right)$ (A.20)
$\geq$ $\mathrm{P}\left(\mathrm{MSS}\subseteq\widehat{\mathrm{MSS}}_{p}\right)-\mathrm{P}\left(\widehat{\mathrm{MSS}}_{p}\nsubseteq\mathrm{MSS}\right).$
In the proof of Lemma A6, we have shown that
$\underset{T\rightarrow\infty}{\lim\sup}\mathrm{P}\left(\widehat{\mathrm{MSS}}_{p}\nsubseteq\mathrm{MSS}\right)\leq p+\varepsilon.$
This, together with (A.12) and (A.20), implies that
$\underset{T\rightarrow\infty}{\lim\inf}\mathrm{P}\left(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p}\right)\geq 1-p-\varepsilon.$ (A.21)
@@
Since $\varepsilon$ is arbitrary, the second claim of the theorem follows from (A.19) and (A.21). $Q.E.D.$
Appendix A.2 Auxiliary Lemmas
@@
This section presents some auxiliary results that are utilized in proving Lemma 1 and Theorem 1 in the previous section. For any $i=1,\ldots,d$, let $\mu_{i}\equiv$ $\mathrm{E}[R_{i}]$, $\mu_{-i}\equiv$ $\mathrm{E}[R_{-i}]$, $\sigma_{i}^{2}\equiv\mathrm{Var}(R_{i})$ and $\Sigma_{-i}\equiv\mathrm{Var}(R_{-i})$. Define $\tilde{\sigma}_{i}^{2}\equiv\sigma_{i}^{2}-\Gamma_{i,-i}\Sigma_{-i}^{-1}\Gamma_{i,-i}^{\top}$, where $\Gamma_{i,-i}\equiv\mathrm{Cov}(R_{i},R_{-i})$.
@@
where $\tilde{M}_{i}\equiv\left|(\ell_{d,i}^{\top}\otimes\mathrm{diag}(s_{i,1}^{-1},s_{i,2}^{-1}))\Omega_{d}^{1/2}\mathcal{N}_{d}\right|_{\infty}$ and $\Omega_{d}=\lim_{T\rightarrow\infty}\Omega_{d,T}$.
For the $t$th observation $R_{t}^{b}$ in the bootstrap sample $\{R_{t}^{b}\}_{t\leq T_{B}}$, let $R_{-i,t}^{b}$ denotes its subvector excluding $R_{i,t}^{b}$. Define $\tilde{R}_{-i,t}^{b}\equiv(1,(R_{-i,t}^{b})^{\top})^{\top}$ and $\varepsilon_{i,t}^{b}\equiv R_{i,t}^{b}-\theta_{i}^{\top}\tilde{R}_{-i,t}^{b}$. For any proper subset $\mathcal{S}$of $\mathcal{M}$, including the empty set, let $cv_{1-p}^{u}(\mathcal{S})$ represent the ($1-p$)th quantile of $\max_{i\notin\mathcal{S}}\tilde{M}_{i}$, and $cv_{1-p,T}^{b}(\mathcal{S})$ represent the ($1-p$)th quantile of $\max_{i\notin\mathcal{S}}M_{i,T}^{b}$. Moreover, we define:
@@
Under Assumptions 1 and 2, we have $cv_{1-p,T}^{b}(\mathcal{M})=O_{p}(1)$ for any $p\in(0,1)$.
@@
Under Assumptions  1 and 2, we have $\lim\inf_{T\rightarrow\infty}\mathrm{P}(\widehat{\mathrm{MSS}}_{p}\subseteq\mathrm{MSS})\geq 1-p$.